OpenStudy (anonymous):

Find the product of 2x4(2x2 + 3x + 4) 2x8 + 3x4 + 4x4 4x6 + 6x5 + 8x4 4x4 + 3x5 + 2x6 3x6 + 4x5 + 5x4

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@paki @Luigi0210 @SithsAndGiggles @pooja195

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

we have to apply the distributive peroperty of multiplication over addition. For example: a*(b+c+d)= a*b+ a*c +a*d

OpenStudy (anonymous):

okay i kinda understand

OpenStudy (michele_laino):

for example: \[\Large \begin{gathered} 2{x^4}\left( {2{x^2} + 3x + 4} \right) = \hfill \\ = \left( {2{x^4} \cdot 2{x^2}} \right) + \left( {2{x^4} \cdot 3x} \right) + \left( {2{x^4} \cdot 4} \right) = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

ohh okay so is it A

OpenStudy (michele_laino):

are you sure?

OpenStudy (anonymous):

well now that i think about it no..

OpenStudy (michele_laino):

hint: \[\Large \begin{gathered} 2{x^4}\left( {2{x^2} + 3x + 4} \right) = \hfill \\ = \left( {2{x^4} \cdot 2{x^2}} \right) + \left( {2{x^4} \cdot 3x} \right) + \left( {2{x^4} \cdot 4} \right) = \hfill \\ = 4{x^6} + ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

its B!!

OpenStudy (michele_laino):

yes! That's right!

OpenStudy (anonymous):

Thanks alottt !! :)

OpenStudy (michele_laino):

thanks! :)

OpenStudy (anonymous):

anytime :)